OpenStudy (anonymous):

A student has 5 nickles and 4 quarters. She picks two randomly. What is the expected total value in cents of the coins?

OpenStudy (anonymous):

@satellite73 @DanielxAK @mathmate I really need to know this for my midterm tomorrow. Even if it says I'm not viewing the question, please answer. I'll check tomorrow morning before I leave for the test.

OpenStudy (anonymous):

@Dido525

OpenStudy (anonymous):

Number of ways to choose the Two coins is \[n C k = 9C2 = 36\] 2 nickels can be picked in \[n C k = 5C3 = 10\] probability = \[\frac{ 10 }{ 36 } =\frac{ 5 }{ 18 }\] 2 quarters can be picked in \[n C k = 4C2 = 6\] Probability = \[\frac{ 6 }{ 36 }= \frac{ 1 }{6 }\] 1 dime and 1 quarter can be picked in \[1-P(quarters)-p(dimes)= \frac{ 5 }{ 9 }\] Expected value:\[P(quarter,dime)(0.30)+P(quarter)(0.50)+P(nickels)(0.10)=value\] =\[(\frac{ 5 }{ 9})(0.30) + (\frac{ 1 }{ 6 })(0.50)+(\frac{ 5 }{ 18 })(0.10) = 27.78=250/9\] in cents.

OpenStudy (anonymous):

@mathsss

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome :) . Hope your midterm was good :) .